### v1

# Quantization-Free Autoregressive Action Transformer

Abstract
Current transformer-based imitation learning approaches introduce discrete action representations and train an autoregressive transformer decoder on the resulting latent code. However, the initial quantization breaks the continuous structure of the action space thereby limiting the capabilities of the generative model. We propose a quantization-free method instead that leverages Generative Infinite-Vocabulary Transformers (GIVT) as a direct, continuous policy parametrization for autoregressive transformers. This simplifies the imitation learning pipeline while achieving state-of-the-art performance on a variety of popular simulated robotics tasks. We enhance our policy roll-outs by carefully studying sampling algorithms, further improving the results.
Machine Learning, ICML
1 Introduction
Generative modeling lies at the heart of modern machine learning, enabling the production of outputs that adhere to user-specified objectives and constraints by navigating high-dimensional spaces of potential outcomes. Fundamentally, this process can be understood as a form of search where naive or exhaustive strategies are computationally intractable due to the size of the space. We consider imitation learning for robotic systems as an extreme example, where the search space has uncountable cardinality and a complex structure arising from nonlinear dynamics, interactions with the environment and the fundamental multimodality in human or robot behavior.
In many real-world robotics applications, the control signals involved are continuous in nature (Muehlebach & D’Andrea, 2017; Zughaibi et al., 2024; Ma et al., 2023). Current state-of-the-art methods for modeling policies in continuous control tasks broadly fall into two categories: autoregressive transformer models (Lee et al., 2024; Shafiullah et al., 2022) and diffusion models (Chi et al., 2023; Wang et al., 2022; Psenka et al., 2023), which will be discussed in the next two paragraphs.
Existing autoregressive policies, on the one hand, sidestep the challenge of learning in a continuous domain by discretizing the actions (Lee et al., 2024; Shafiullah et al., 2022). This discretization can introduce several drawbacks: It discards the inherent structure of the continuous space, increases complexity by adding a separate quantization step, and may limit expressiveness or accuracy when fine-grained control is required. Furthermore, the popular approach of learning a VQ-VAE (Van Den Oord et al., 2017) of the action space involves non-differentiable operations and hence requires advanced tricks during representation learning to side-step optimization difficulties and instabilities (Łańcucki et al., 2020; Kolesnikov et al., 2022; Huh et al., 2023; Mentzer et al., 2024). Hence, our goal is to design autoregressive transformers that preserve the continuous nature of the action space, leveraging the smooth and potentially multimodal behavior patterns that experts exhibit.
Figure 1: We present the diversity of behaviors generated by Q-FAT across three different environments. The first three images illustrate the initial task execution in the Franka Kitchen environment (Gupta et al., 2019), where the robot begins by either moving the kettle, opening the microwave, or interacting with the bottom burner. The third and fourth images showcase the UR3 BlockPush environment, demonstrating how the robotic manipulator initiates the task by pushing either the blue or red block. Finally, the last image depicts the PushT environment (Chi et al., 2023), where Q-FAT successfully captures the multimodal nature of the task, producing smooth and diverse motion trajectories.
Diffusion policies, on the other hand, are adept at capturing continuous action distributions, but face two critical challenges in control tasks. First, real-time applications like robotics demand fast, efficient inference, yet diffusion models rely on iterative sampling, requiring dozens of forward passes to generate actions. This makes them orders-of-magnitude slower than autoregressive or feedforward policies (Lee et al., 2024). Second, the downstream tasks of imitation learning is often reinforcement learning (Schaal, 1996; Peters & Schaal, 2008; Zhu et al., 2018; Hu et al., 2024), which typically requires exact action likelihoods for tasks like exploration and off-policy corrections (Sutton et al., 1999; Abdolmaleki et al., 2018; Levine et al., 2020; Wu et al., 2019). Diffusion models, however, generate samples by integrating (stochastic) differential equations (Song et al., 2023, 2020), making likelihood estimation computationally expensive.
To address these limitations, we propose the use of a Generative Infinite Vocabulary Transformer (GIVT) (Tschannen et al., 2024) that models the policy as a Gaussian Mixture Model (GMM) on top of a decoder-only transformer architecture (Vaswani, 2017; Radford et al., 2018). This approach enables modeling continuous actions as “tokens“ by directly parameterizing the components of the GMM in each time step, avoiding the need for discretization and thus preserving the inherent geometry of the action space. This choice allows us to exploit the benefits of large-scale sequence modeling, such as autoregressive policy generation and access to explicit likelihood estimates, while avoiding the pitfalls of artificially segmenting continuous control signals into discrete bins. In addition, we discuss two different sampling strategies that reduce variability in the generated trajectories.
In summary, our work:
  • Eliminates the need for action quantization in autoregressive policy parameterization, simplifying behavioral cloning pipelines while retaining the ability to directly evaluate action likelihoods.
  • Proposes two sampling algorithms to reduce the variance of the generated trajectories, including a global mode-finding algorithm.
  • Demonstrates state-of-the-art performance on standard benchmarks, both for conditional and unconditional policy generation using proprioceptive and image-based states, confirming the effectiveness of our method for multimodal policy modeling.
The implementation is available at https://github.com/ziyadsheeba/qfat.
Figure 2: Overview of Q-FAT. A sequence of $h^{s}$ previous states and $h^{g}$ goal states are projected into the Transformer’s embedding dimension using a linear layer. The Transformer then predicts the action distribution by predicting the GMM means, variances and mixture probabilities.
2 Sequence to Action Prediction
We consider the following standard setting for imitation learning of simulated robotics tasks. An agent equipped with continuous controls $\mathcal{A}$ evolves in a continuous state space $\mathcal{S}$. In behavioral cloning, we assume that a so-called behavior policy $\pi_{b}$ collected a dataset $D=\{\tau_{i}\}_{i=0}^{N-1}$ of $N$ trajectories in the environment, where each trajectory $\tau_{i}=\{s^{i}_{0},a^{i}_{0},\dots,a^{i}_{n^{i}-1},s^{i}_{n^{i}}\}$ is a sequence of states $s^{i}_{t}\in\mathcal{S}$ and actions $a^{i}_{t}\in\mathcal{A}$, with $n^{i}$ denoting the length of trajectory $i$. Note that we do not impose a Markovian assumption on the behavior policy: in general $a^{i}_{t}$ might depend on all previous states $s^{i}_{<t}$.
Our objective is to learn a sequence-to-action predictor, or imitating policy $\pi_{\theta}$. We treat the objective as a supervised learning problem, where the parameter $\theta$ of our policy is optimized to minimize the negative log-likelihood with respect to $\pi_{b}$, $\mathbb{E}_{\tau\sim\pi_{b}}[-\log(\pi_{\theta})(\tau)]$. To do so, we minimize:
$\mathcal{L}(\theta)=-\frac{1}{N}\sum_{i=0}^{N-1}\sum_{t=0}^{n^{i}-1}\log\pi_{\theta}(a^{i}_{t}\mid s^{i}_{<t})\,.$ (1)
Note that a standard result in variational inference (Hoffman et al., 2013) shows that this objective is equivalent to minimizing the Kullback-Leibler (KL) divergence between the behavioral policy $\pi_{b}$ and the learned policy $\pi_{\theta}$, implying that the learnt policy $\pi_{\theta}$ should match, or clone, the behavior of $\pi_{b}$. However, the choice of parametrization is crucial to guarantee a good approximation of the behavior policy.
3 The Q-FAT Algorithm
Current transformer-based policy parametrization approaches (Lee et al., 2024; Shafiullah et al., 2022; Mete et al., 2024) consist of first discretizing the action space, e.g. using a VQ-VAE (Van Den Oord et al., 2017; Zeghidour et al., 2021), to comply with the standard transformer architecture requiring a discrete vocabulary. Inspired by the success of GIVT  (Tschannen et al., 2024) in the image generation domain, our model Quantization-Free Autoregressive Action Transformer (Q-FAT) adapts GIVT to the continuous action behavioral cloning setting, modeling the action distribution at each step by predicting the parameters of a multivariate Gaussian mixture model (GMM). Namely, Q-FAT models the probability of an action $a_{t}\in\mathbb{R}^{m}$ at step $t$ as
$\pi_{\theta}\bigl{(}a_{t}\,\big{|}\;s_{t-1},\dots,s_{t-h^{s}})\!=\!\sum_{i=1}^{k}\pi_{i}^{t}\,\mathcal{N}\!\Bigl{(}a_{t}\;\big{|}\;\boldsymbol{\mu}_{i}^{t},\boldsymbol{\sigma}_{i}^{t}\Bigr{)}\,,\quad$ (2)
where $\pi_{i}^{t}$, $\boldsymbol{\mu}_{i}^{t}$, and $\boldsymbol{\mathbf{\sigma}_{i}^{t}}$ are produced by the transformer decoder (Vaswani, 2017; Radford et al., 2018) at step $t$, conditioned on a history of previous states $\{s_{i}\}_{i=t-h^{s}}^{t-1}$ with $s_{i}\in\mathbb{R}^{d}$. Using a state history of length $h^{s}$ instead of only using the most recent state accounts for the potential non-Markovian structure in the demonstration dataset. Rather than predicting the full covariance $\mathbf{\Sigma}^{t}_{i}$ per mixture component, we assume a diagonal Gaussian distribution and only predict the diagonal entries of the covariance matrix $\boldsymbol{\sigma}^{t}_{i}$. While this simplifying assumption enables efficient likelihood computations, it does not impose independence across action dimensions in general.
3.1 Goal Conditioning
For some tasks such as tracking, a near-future goal can be appended to the input to condition the resulting policy. The conditioning is formalized by defining a set of goal states $\mathcal{G}\subseteq\mathcal{S}$. We then learn a goal-conditioned policy $\pi_{\theta}\bigl{(}a_{t}\,\big{|}\;s_{t-1},\dots,s_{t-h^{s}},g\bigr{)}$ where $g\subset G$. In practice, we define $g$ as a sequence of future states to be reached, i.e., $g=\{s^{g}_{i}\}_{i=t}^{t+h^{g}}$, where the goal states $s^{g}_{i}$ are typically generated from a high-level planner that plans instrumental goal states as done in Gupta et al. (2019). In our experiments, we simulate goal states from the demonstration dataset.
3.2 Action Sampling from a $k$-GMM
Q-FAT samples the next action from the conditional distribution, $a_{t}\sim\pi_{\theta}(\cdot|s_{t-1},\dots,s_{t-h^{s}},g)$, which is a GMM with $k$ components. In practice, the model can be trained to output a short sequence of actions that are then executed in an open-loop fashion. This is done to ensure action consistency and avoid over-fitting to idle actions (Chi et al., 2023). For simplicity, we describe only the next-action generation setting.
One can directly sample from the GMM that parametrizes the policy. For the vanilla GMM sampling, the procedure is:
  1. 1.
    Sample a mixture index $k$ from the categorical distribution defined by the probabilities $\{\pi_{i}\}_{i=1}^{k}$.
  2. 2.
    Sample an action $\mathbf{a}$ from the corresponding Gaussian component $\mathcal{N}\bigl{(}\boldsymbol{\mu}_{i},\boldsymbol{\sigma}^{2}_{i}\bigr{)}$.
While this approach directly samples from the learned policy distribution, it can degrade performance in practice when the dataset is noisy. In that case, the GMM may overfit the data distribution and produce noisy actions, resulting in jitter or unstable trajectories. We discuss below how Q-FAT can be easily equipped with stabilizing sampling techniques that directly exploit the fitted GMM distribution.
3.3 Stabilizing output trajectories
When controlling real-world robotics systems, it is important to generate action sequences without jitter, as high frequency inputs may damage actuators and excite undesirable structural modes (Astrom, 1995). Thanks to the richness of the GMM representation of the policy, we can directly design stabilization patches.
Figure 3: Potential detrimental effect of down-scaling the variance of the GMM mixture components when $k$ is misspecified. A uni-modal Gaussian distribution (gray) is approximated with two components (orange and green). The resulting density function (blue) fits the data distribution perfectly. Down-scaling the variance to $0.1$ (red) induces irreducible variance proportional to the distance between the modes.
Down-scaling the Mixture Variances
A common heuristic to mitigate noisy samples is to adjust the sampling temperature, which in our GMM parametrization is equivalent to down-scaling the variance in each Gaussian component (Mandlekar et al., 2021; Hoffman et al., 2020; Tschannen et al., 2024, 2025). Concretely, one replaces each covariance $\boldsymbol{\sigma}_{i}$ by $\alpha\,\boldsymbol{\sigma}_{i}$, for some small $\alpha\in(0,1)$—sometimes by several orders of magnitude ($\alpha\ll 1$). This compresses each component, forcing sampled actions to concentrate more tightly around each component’s mean. However, merely shrinking the variances does not address the large-scale spread when mixture components are far apart. The inter-component variance remains, and noisy samples may still arise from small but non-negligible transitions among disparate components. This phenomenon is exacerbated when the choice of the number of components $k$ is misspecified (see Figure 3 and further discussion in Appendix A). Furthermore, it has been shown by Carreira-Perpinan (2000) that GMMs in dimension $d\geq 2$ may have a mode that does not directly align with the mean of any component. In such a situation, the down-scaling technique is likely to result in an insidious loss of the extra mode (see Figure 4).
Mode-tracking via the Mean-Shift algorithm We explored a principled way to reduce the variance of sampling from multimodal distributions by explicitly identifying the major modes and then sampling from those modes with an appropriate probability while controlling variance. An effective way to find the modes in a GMM is via the mean-shift algorithm (Carreira-Perpinan, 2000), which is an efficient fixed-point iteration algorithm (see Appendix B for further details). We propose a multinomial distribution over the modes [^1] The number of modes $J$ is hard to control theoretically (Carreira-Perpinan, 2000). $\{m_{1},...m_{J}\}$ that respects the shape of the density:
$\Pr(m_{j})$ $=\frac{w_{j}}{\sum_{j^{\prime}}w_{j^{\prime}}}\,,$ (3)
where $\text{where }w_{j}$ $=p_{\text{GMM}}\bigl{(}m_{j}\bigr{)}\times\bigl{|}-H_{j}\bigr{|}^{-\frac{1}{2}}$ (4)
and $\text{and }H_{j}$ $=\nabla^{2}\log p_{\text{GMM}}(m_{j})\,.$ (5)
Note that each $H_{j}$ admits a closed-form expression and can be computed in $O(km^{2})$ operations. One can then further inject noise to the sampled mode using a Gaussian noise with a fixed variance or use the Hessian around the mode to account for the curvature of the distribution to determine the variance(see more details on this process in Appendix B). We emphasize that this mode-tracking algorithm can be efficiently implemented on GPUs using common deep learning libraries and introduces a minor overhead compared to the transformer forward pass.
[…]
(a) Demonstration Dataset
(b) GMM Sampling
(c) $\sigma$-scaled Sampling
(d) Mode Sampling
Figure 5: We demonstrate the effect of multiple sampling techniques from a $8$-mixture Q-FAT policy trained on the toy multi-route environment (Shafiullah et al., 2022). (a) Visualization of the raw dataset, where the agent starts from the bottom-left corner and follows either the blue or the red path, with the goal of landing inside of the green target region on the top-right corner. Both paths are equally likely in the dataset. (b) Sampling directly from the GMM results in noisy samples because the model also captures and samples from the noise present in the dataset. (c) In $\sigma$-scaled sampling, the variances of the GMM components are scaled by a factor of $10^{-8}$. Despite reducing the variance per component, variance scaling does not decrease for the variance introduced due to the distance between the mixture components. (d) Mode sampling suppresses most of the noise in the demonstrations while preserving the multimodality of the original dataset.
4 Experiments
We evaluate the performance of Q-FAT through extensive experiments across four representative simulated robotics environments: PushT (Chi et al., 2023), Kitchen (Gupta et al., 2019), UR3 BlockPush (Florence et al., 2022) and Multimodal Ant (Lee et al., 2024). These environments cover the most common control signals in practice, namely position and velocity of robot’s end effector and robot joint angles. A detailed description of each environment is provided in Appendix C. Our experiments are designed to address the following key research questions:
  1. 1.
    Is Q-FAT competitive with state-of-the-art methods in both conditional and unconditional behavior imitation?
  2. 2.
    Can Q-FAT effectively capture the diversity of demonstration datasets while avoiding mode collapse?
  3. 3.
    What is the effect of different sampling techniques on the performance of Q-FAT?
  4. 4.
    How sensitive is Q-FAT’s performance to the choice of the mixture components $k$ in the learned policy?
We find that Q-FAT achieves state-of-the-art performance across the majority of conditional and unconditional tasks, achieving state-of-the-art balance between the quality and diversity of the generated actions. We report the results of Q-FAT using the variance reduced sampling (Section 3.2) with a variance scaling of $10^{-6}$ and we discuss the implications of this choice in Section 4.3. A summary of our results is presented in Table 1.
4.1 Baselines and Performance Metrics
— Unconditional Tasks —
Environment Metric BeT DiffPolicy-C DiffPolicy-T VQ-BeT Q-FAT
PushT Final IoU (/1$\cdot$/1⋅ / 1) 0.39 0.73 0.74 0.78 0.80
Image PushT 0.01 0.66 0.45 0.68 0.69
Kitchen Goals (/4$\cdot$/4⋅ / 4) 3.07 2.62 3.44 3.66 3.84
Multimodal Ant 2.73 3.12 2.90 3.22 3.30
UR3 BlockPush Goals (/2$\cdot$/2⋅ / 2) 1.59 1.83 1.82 1.84 1.96
— Conditional Tasks —
Environment Metric C-BeT C-BESO CFG-BESO VQ-BeT Q-FAT
Conditional Kitchen Goals (/4$\cdot$/4⋅ / 4) 0.15 3.75 3.47 3.78 3.78
Conditional UR3 BlockPush Goals (/2$\cdot$/2⋅ / 2) 0.00 1.14 0.92 1.94 1.96
Table 1: Performance comparison between unconditional and conditional behaviors on the task success rates. We report the performance of Q-FAT with variance down-scaling with a factor of $10^{-6}$.
Baselines
We compare Q-FAT against a range of state-of-the-art baselines. We restrict our focus to comparing against diffusion and transformer-based policy learning methods. For unconditional tasks, the baselines include: (1) BeT (Shafiullah et al., 2022), which performs action discretization using k-means clustering and models the action distribution with a transformer decoder; (2) VQ-BeT (Lee et al., 2024), an extension of BeT that leverages vector quantization via a hierarchical Variational Autoencoder (VQ-VAE) (Zeghidour et al., 2021) for action discretization; and (3) two variants of Diffusion Policy (Chi et al., 2023), namely a convolutional-based implementation and a transformer-based implementation. For goal-conditional tasks, we extend the comparison to include conditional versions of BeT and VQ-BeT, as well as two additional baselines: BESO (Reuss et al., 2023), a denoising diffusion-based method with a conditional variant (C-BESO) and a classifier-free guided variant (CFG-BESO).
4.2 Quality and Diversity of Q-FAT’s Actions
Figure 6: Overview of Q-FAT’s behavioural entropy on unconditional behavior generation compared to baselines.
Evaluation Metrics
We assess performance using task success rate and induced behavioral entropy. Task success is measured differently per environment: in Kitchen, Multimodal Ant and UR3 BlockPush, the task success rate corresponds to the number of goals achieved per roll-out, while in PushT, it is quantified using the final Intersection over Union (IoU) between the object’s position and the target area. Behavioral diversity is measured by computing task completion sequence entropy (length 4 for Kitchen and Multimodal Ant, length 2 for UR3 BlockPush), reflecting variability in task completion. Baseline performance follows metrics from Lee et al. (2024), and results are reported as the average success rate of the best checkpoint over 1000 episodes, including behavioral entropy where applicable.
Comparison to Baselines
Q-FAT achieves state-of-the-art performance on the success metrics, as shown in Table 1. The performance gap observed with BeT can be attributed to the limitations of k-means clustering in high-dimensional spaces, which may hinder its ability to effectively capture complex action distributions. This is particularly exacerbated due to BeT attempting to discretize the whole unconditional action space, and not the conditional one. In the case of VQ-BeT, the sequential nature of first predicting a discrete action token followed by an action correction step introduces two potential sources of error: inaccuracies in token prediction and errors in the subsequent correction mechanism. This places significant reliance on the vector quantization network, which is challenging to optimize due to the need to discretize the entire action space and the two-stage training process required—first for the quantization network and then for the policy. For diffusion-based policies, the absence of an explicit likelihood function makes it difficult to control the variance of sampled actions, particularly in the presence of noisy data. This can be particularly problematic in fine-grained control tasks, where precise action execution is crucial for success. Consequently, while these approaches demonstrate strong performance in certain scenarios, they may struggle in tasks requiring high accuracy and reliability under noisy data. As for assessing the behavioral diversity, Q-FAT largely outperforms baselines on the behavioral entropy in the Kitchen and the Multimodal Ant environments, while marginally under-performing in UR3 BlockPush compared to BeT, VQ-BeT and the transformer-based Diffusion Policy (Figure  6).
4.3 Effects of Sampling Techniques
We assessed the effect of the different sampling techniques discussed in Section 3.2. We found that down-scaling the component variances had a 7% and a 6.5% increase in the success performance metrics in the Kitchen environment and PushT, respectively, compared to the GMM sampling, while only seeing a marginal 2% drop in the diversity of the generated behaviors in the Kitchen environment. The modes sampling matched the performance of the variance down-scaling on the evaluation metrics. This could be attributed to the fact that the learned mixture components are sufficiently distant in the environments we explored, thus the effects discussed in Section 3.2 are not pronounced. For UR3 BlockPush, the sampling methods performed similarly on the performance metrics. In the Multimodal Ant environment, we observed a substantial 65% decrease in performance when using variance downscaling and mode sampling at $k=2$, compared to direct sampling from the GMM. Notably, this significant performance drop was only observed at $k=2$; at higher values of $k$, performance differences between all sampling methods became negligible. We expect the performance gap between the variance reduction sampling and the modes sampling to become more severe in higher-dimensional action spaces, but leave this exploration for future work.
Despite the quantitative metrics not showing significant differences between the variance down-scaling method and the mode sampling, we visualize the trajectories produced by both sampling techniques on a toy multi-route environment introduced in Shafiullah et al. (2022) and see the effects of the irreducible variance in Figure 5. We further display a similar effect in the PushT environment in Figure 7.
Figure 7: We visualize 400 trajectories generated from Q-FAT with 16 mixtures on the PushT environment with different sampling methods. One can see that mode sampling reduces the variance and does not produce trajectory artifacts.
4.4 Effect of the Number of Mixtures
We systematically evaluated the effect of varying the number of mixture components $k$ in Q-FAT, where $k$ controls the number of modes in the conditional action distribution for a given state context. Specifically, we considered $k\in\{2,4,8,16\}$ in two benchmark domains, the Kitchen and UR3 BlockPush environments. Our results indicate that Q-FAT’s success performance is highly robust across these values of $k$, with negligible degradation in performance. In the Kitchen environment, we observed an 18% increase in behavioral entropy when $k$ was increased from 2 to 4, after which entropy remained approximately constant as $k$ continued to increase. In contrast, in the UR3 BlockPush environment, the behavioral entropy remained stable regardless of $k$.
The key to achieving this robustness is careful initialization. In particular, we found it important to initialize each mixture component’s mean to be far from the others. Concretely, we placed these means on a hyper-cube spanning the range $[-1,1]$ per dimension, maximizing their mutual separation, and assigned each component a unit variance. The range $[-1,1]$ is a reasonable choice, as actions are typically bounded in practice and normalized in our experiments.
To investigate how the mixture components were utilized during inference, we tracked the average number of active components over the course of training (see Figure 8). We observed that the model effectively prunes redundant mixture components as training progresses, with the pruning effect becoming more pronounced when more components are initially employed. Interestingly, in the Kitchen environment, which has a nine dimensional action space, less pruning occurs as $k$ increases compared to the UR3 environment, whose action space is only two dimensional.
A plausible explanation for this discrepancy lies in the dimensionality of the action space. In higher dimensions, each Gaussian component covers a proportionally smaller volume of the action space. Consequently, the model requires more components to adequately represent the conditional action distribution, and therefore prunes fewer mixture components. In lower-dimensional environments, by contrast, each component can cover a larger fraction of the data, which makes aggressive pruning more viable.
We further plotted the distribution over the number of active modes for the Kitchen and the UR3 BlockPush environments. We observed that approximately 70% of the course of a trajectory is uni-modal (Figure 9). This shows that multi-modality is required only a minority of the time during a rollout. This observation could be useful for future work to exploit this uni-modality computationally.
Figure 8: We observe that Q-FAT has a pruning effect, reducing the number of active mixtures at inference as training proceeds. We define the active number of mixtures at each training step as the count of mixture components whose probability is at least $10^{-2}$ times the maximum mixture probability at each inference step. We averaged the number of active mixtures over the rollout horizon for each environment and display the ratio of active mixtures for each $k$ value.
Figure 9: We analyzed the mode distribution over 100 rollouts for trained models with $k=4$ in the Kitchen and UR3 environments. Our findings indicate that in both environments, approximately 70% of the trajectory maintains only a single active mode, suggesting that the policy remains largely unimodal throughout each rollout.
4.5 Training Details
For our experiments, we used the minGPT (Brown et al., 2020) backbone as the decoder-only transformer implementation. We added the hyperparameters used for each of the environments in Table 2.
During training, we observed instances where both validation and training losses decreased, yet the environment rewards declined. We attribute this phenomenon to causal confusion (De Haan et al., 2019), where the model overfitslations in the recent sequence of states. Instead of learning true causal relationships, it relies on temporally correlated but non-causal patterns to predict future actions, leading to low rewards. To mitigate this issue, we introduced history masking, wherein ein whichment in the context (except for the current state) is masked with a certain probability. We applied this technique in both the Kitchen and Multimodal Ant environments (with probabilities 0.7 and 0.3 respectively), which led to an improved correlation between environment rewards and validation loss.
For the Image PushT, we fine-tuned a ResNet18 encoder (He et al., 2016) by extracting the feature maps from the first four layers in order to construct low level features (Parisi et al., 2022), which are then spatially average-pooled and fed into the transformer with a dimensionality of 1024. We further applied data augmentation, namely color jitter, random cropping, and random gray-scaling of the image with a probability of 0.5. We found this to be important to preserve the invariance of the pre-trained ResNet and not overfit to our dataset.
In the goal conditional tasks, we used a sequence of future states that has the same length as the state history. At inference time, we randomly sample a trajectory from the dataset and track it, computing the reward only if the agent achieved the goal in the same order that was achieved by the reference trajectory.
5 Related Work
Offline Learning for Decision Making
Learning to act from offline data has received significant attention over the years due to its vast potential in practical applications. This line of research can be divided into two categories: offline reinforcement learning (Lange et al., 2012; Levine et al., 2020; Kumar et al., 2020; Chen et al., 2021) and imitation learning (Pari et al., 2021; Peng et al., 2018, 2021; Ho & Ermon, 2016). In offline reinforcement learning, on the one hand, the goal is to learn a policy from demonstrations with suboptimal behaviors, which necessitates the access to rewards in the demonstration dataset. Imitation learning, on the other hand, assumes that the trajectories are collected from near optimal behaviors, avoiding access to reward signals in the collected datasets. This increases the attractiveness of imitation learning in practical setups, since reward signals are hard to define (Devidze et al., 2021; Kumar et al., 2022). Q-FAT falls under the framework of behavioral cloning, which is a form of imitation learning where the policy learning is treated as a supervised learning problem.
Generative Models for Cloning Behavior
Complex generative models have been employed to capture the full data diversity due to the complex and multimodal behavior in human and robot collected datasets. Earlier work has explored the use of Gaussian processes (Shon et al., 2005), energy-based models (Florence et al., 2022) and generative adversarial networks (Kuefler et al., 2017) as policy parametrizations. More recently, more attention has been directed towards more powerful models, such as transformers (Shafiullah et al., 2022; Lee et al., 2024; Mete et al., 2024; Wu et al., 2023) and diffusion models (Chi et al., 2023; Wang et al., 2022; Psenka et al., 2023; Pearce et al., 2023). Unlike other transformer-based policies, Q-FAT uses a continuous parametrization of the output by predicting a GMM. Note that this idea has been explored before using LSTM backbones (Mandlekar et al., 2021; Shafiullah et al., 2022), however, it has been shown to produce poor performance (Shafiullah et al., 2022; Lee et al., 2024), which we attribute to the instability of training LSTM networks and their limited capacity to model long sequences (Hochreiter, 1997).
6 Discussion and Future Work
In this work, we introduced Q-FAT, a quantization-free action transformer that achieves state-of-the-art performance across various simulated robotics environments, opening new research directions in both behavioral modeling and reinforcement learning. Compared to prior work  (Lee et al., 2024), Q-FAT overcomes the challenging non-differentiable step in action discretization while maintaining or improving performance.
By integrating Q-FAT into end-to-end policy learning pipelines, future research could evaluate its influence on sample efficiency and final task performance. Another promising area involves incorporating Bayesian priors into Q-FAT’s action distribution estimates, which could facilitate active exploration in complex, high-dimensional settings. For fine manipulation tasks, researchers could further investigate coarse-to-fine sampling strategies based on Gaussian mixture model representations, potentially improving both exploration breadth and control. Finally, extending Q-FAT to non-Euclidean action spaces, such as those with Riemannian geometry, may enable more accurate and natural representations for tasks like legged locomotion or dexterous manipulation.
Impact Statement
Our work tackles simulated continuous control tasks whose downstream consequences are important for robotics. The probabilistic nature of our method may impact the safety of the generated behavior and should be further studied and tested. However, unlike prior work, our approach enables uncertainty estimates directly in the action space, which have the potential to improve safety by providing uncertainty quantification.
Acknowledgment
We would like to thank Youssef Saied for the valuable discussions throughout the project and Nicolas Nguyen for reviewing the document. C.V. and Z.S. are funded by the German Research Foundation (DFG) under both the project 468806714 of the Emmy Noether Programme and under Germany’s Excellence Strategy – EXC number 2064/1 – Project number 390727645. M.M. is funded by the German Research Foundation (DFG) under the project 456587626 of the Emmy Noether Programme. We also thank the international Max Planck Research School for Intelligent Systems (IMPRS-IS) for the support.
Appendix A GMM Properties
GMM Moments
Let $\mathbf{X}$ be a random variable from a $k$-mixture GMM distribution in $\mathbb{R}^{d}$. The probability density function can then be written as:
$p(\mathbf{x})=\sum_{i=1}^{k}\pi_{i}\mathcal{N}(\mathbf{x}\mid\boldsymbol{\mu}_{i}\,,\boldsymbol{\Sigma}_{i})\,,$ (6)
where $\mathcal{N}(\mathbf{x}\mid\boldsymbol{\mu}_{i},\boldsymbol{\Sigma}_{i})$ is the multivariate Gaussian distribution for the $i$-th component. The mean and the covariance of $\mathbf{X}$ are then given by:
$\mathbb{E}[\mathbf{X}]$ $=\sum_{i=1}^{k}\pi_{i}\boldsymbol{\mu}_{i},$ (7)
$\mathrm{Cov}[\mathbf{X}]$ $=\sum_{i=1}^{k}\pi_{i}\left(\boldsymbol{\Sigma}_{i}+(\boldsymbol{\mu}_{i}-\mathbb{E}[\mathbf{X}])(\boldsymbol{\mu}_{i}-\mathbb{E}[\mathbf{X}])^{\top}\right)\,,$ (8)
where $\pi_{i}$ are the mixing coefficients, and $\boldsymbol{\mu}_{i}$, $\mathbf{\Sigma}_{i}$ are the mean and covariance parameters for the $i$-th Gaussian component. Note that while the mean of the GMM is a weighted sum of the component means, the covariance is not. The total covariance reflects both the inherent variability within each component and the variability introduced by the differences in component means.
Appendix B GMM Mode Finding and Sampling
Mode Finding
To find the modes of the GMM, we compute the gradient of the probability density function with respect to $\mathbf{x}$ and equate it to zero:
$\nabla p(\mathbf{x})=\sum_{i=1}^{k}\pi_{i}\nabla\mathcal{N}(\mathbf{x}\mid\boldsymbol{\mu}_{i},\boldsymbol{\Sigma}_{i})=\mathbf{0}\,,$ (9)
where the gradient of the multivariate Gaussian distribution is given by
$\nabla\mathcal{N}(\mathbf{x}\mid\boldsymbol{\mu}_{i},\boldsymbol{\Sigma}_{i})=\mathcal{N}(\mathbf{x}\mid\boldsymbol{\mu}_{i},\boldsymbol{\Sigma}_{i})\boldsymbol{\Sigma}_{i}^{-1}(\boldsymbol{\mu}_{i}-\mathbf{x})\,.$ (10)
We obtain therefore:
$\sum_{i=1}^{k}\pi_{i}\mathcal{N}(\mathbf{x}\mid\boldsymbol{\mu}_{i},\boldsymbol{\Sigma}_{i})\boldsymbol{\Sigma}_{i}^{-1}(\boldsymbol{\mu}_{i}-\mathbf{x})=\mathbf{0}.$ (11)
We identify the modes of the GMM by computing all critical points $\mathbf{x}^{*}$ that satisfy (11). Moreover, to ensure that a critical point $\mathbf{x}^{*}$ is indeed a mode, the Hessian of the density function at that point must be negative definite. The Hessian of the GMM density is given by
$\nabla^{2}p(\mathbf{x})=\sum_{i=1}^{k}\pi_{i}\nabla^{2}\mathcal{N}(\mathbf{x}\mid\boldsymbol{\mu}_{i},\boldsymbol{\Sigma}_{i}),$ (12)
where the Hessian of the multivariate Gaussian distribution is:
$\nabla^{2}\mathcal{N}(\mathbf{x}\mid\boldsymbol{\mu}_{i},\boldsymbol{\Sigma}_{i})=\mathcal{N}(\mathbf{x}\mid\boldsymbol{\mu}_{i},\boldsymbol{\Sigma}_{i})\left[\boldsymbol{\Sigma}_{i}^{-1}(\boldsymbol{\mu}_{i}-\mathbf{x})(\boldsymbol{\mu}_{i}-\mathbf{x})^{\top}\boldsymbol{\Sigma}_{i}^{-1}-\boldsymbol{\Sigma}_{i}^{-1}\right]\,.$ (13)
Therefore, the Hessian of the GMM density becomes
$\nabla^{2}p(\mathbf{x})=\sum_{i=1}^{k}\pi_{i}\mathcal{N}(\mathbf{x}\mid\boldsymbol{\mu}_{i},\boldsymbol{\Sigma}_{i})\left[\boldsymbol{\Sigma}_{i}^{-1}(\boldsymbol{\mu}_{i}-\mathbf{x})(\boldsymbol{\mu}_{i}-\mathbf{x})^{\top}\boldsymbol{\Sigma}_{i}^{-1}-\boldsymbol{\Sigma}_{i}^{-1}\right]\,,$ (14)
and a critical point $\mathbf{x}^{*}$ is a mode if: $\nabla p(\mathbf{x}^{*})=\mathbf{0}\text{ and }\nabla^{2}p(\mathbf{x}^{*})\text{ is negative definite.}$.
While GMMs are often parameterized by a finite number of components, the resulting distribution can exhibit a richer structure, particularly in higher-dimensional spaces. Specifically, in dimensions greater than two ($d>2$), a GMM can possess more modes (local maxima) than the number of its constituent components (Carreira-Perpinan, 2000). This phenomenon arises due to the interplay between the covariance structures and the relative positions of the Gaussian components. In higher dimensions, the overlapping regions of different components can create multiple peaks in the probability density function that cannot be attributed to individual components.
Mean-Shift
We start by solving (11) for $\mathbf{x}$ as follows:
$\mathbf{x}\;=\;\Biggl{(}\sum_{i=1}^{k}\pi_{i}\,\mathcal{N}\bigl{(}\mathbf{x}\,\big{|}\,\boldsymbol{\mu}_{i},\boldsymbol{\Sigma}_{i}\bigr{)}\,\boldsymbol{\Sigma}_{i}^{-1}\Biggr{)}^{-1}\;\sum_{i=1}^{k}\pi_{i}\,\mathcal{N}\bigl{(}\mathbf{x}\,\big{|}\,\boldsymbol{\mu}_{i},\boldsymbol{\Sigma}_{i}\bigr{)}\,\boldsymbol{\Sigma}_{i}^{-1}\,\boldsymbol{\mu}_{i}\,.$ (15)
To obtain the mean-shift update in practice, one can treat the solution to (15) as a fixed-point iteration. More precisely, we define the operator $T:\mathbb{R}^{d}\to\mathbb{R}^{d}$ by
$T(\mathbf{x})=\Biggl{(}\sum_{i=1}^{k}\pi_{i}\,\mathcal{N}\bigl{(}\mathbf{x}\,\big{|}\,\boldsymbol{\mu}_{i},\boldsymbol{\Sigma}_{i}\bigr{)}\,\boldsymbol{\Sigma}_{i}^{-1}\Biggr{)}^{-1}\;\sum_{i=1}^{k}\pi_{i}\,\mathcal{N}\bigl{(}\mathbf{x}\,\big{|}\,\boldsymbol{\mu}_{i},\boldsymbol{\Sigma}_{i}\bigr{)}\,\boldsymbol{\Sigma}_{i}^{-1}\,\boldsymbol{\mu}_{i}.$ (16)
Then, starting from an initial guess $\mathbf{x}^{(0)}$, the mean-shift procedure updates the estimate of $\mathbf{x}$ via
$\mathbf{x}^{(t+1)}\;=\;T\bigl{(}\mathbf{x}^{(t)}\bigr{)}\,.$
The fixed-point iteration has the following interpretation: each iteration shifts the current iterate after $t$ iterations closer to the modes (high-density regions) of the underlying density defined by the mixture of Gaussians. In practice, it has been shown that initializing $\mathbf{x}^{(0)}$ with the component means $\boldsymbol{\mu}_{i}$ results in fast convergence (Carreira-Perpinan, 2000), which we also observe in our experiments. To account for the fact that the number of modes could exceed the number mixture components, we add extra initialization points from the convex-hull of the component centroids capturing the most important modes.
Mean-Shift with Diagonal-Covariance Components
An important special case arises when the covariance matrices $\boldsymbol{\Sigma}_{i}$ are diagonal. In that scenario, let
$\boldsymbol{\Sigma}_{i}\;=\;\mathrm{diag}\bigl{(}\sigma_{i,1}^{2},\dots,\sigma_{i,d}^{2}\bigr{)}\,.$
Since each $\boldsymbol{\Sigma}_{i}^{-1}$ is also diagonal (with entries $\frac{1}{\sigma_{i,j}^{2}}$ along the diagonal), the vector and matrix operations inside $T(\mathbf{x})$ decouple across dimensions. Hence the inverse of $\sum_{i=1}^{k}\pi_{i}\,\mathcal{N}(\mathbf{x}\mid\boldsymbol{\mu}_{i},\boldsymbol{\Sigma}_{i})\,\boldsymbol{\Sigma}_{i}^{-1}$ reduces to a diagonal matrix whose components can be computed efficiently. As a result, each coordinate of the updated point can be determined independently, making the mean-shift iteration particularly fast on modern hardware.
Modes Sampling
To sample the modes proportional to their coverage of the data support, we approximate the integral of the GMM density in the local neighborhood of each mode using a Laplace approximation and use the value of the integral around each mode to compute the mode’s relative weight. We summarize the full mode-finding and sampling procedure in Algorithm 1.
Algorithm 1 Mode Extraction and Sampling
1:  Inputs: GMM parameters $\{\pi_{i},\boldsymbol{\mu_{i}},\boldsymbol{\Sigma_{i}}\}_{i=1}^{k}$; tolerance $\epsilon$; rejection threshold $\theta$; max iterations max_it; number of initializations n_init;
2:  Initialize Mode Set: $\mathcal{M}\leftarrow\emptyset$
[…]
4:  for $x\in\mathcal{X}$ do
5:     Iteration counter $i\leftarrow 0$
6:     Define the operator $T(.)$ as in Equation 16
7:     repeat
8:        Update $\mathbf{x}$ via fixed-point iteration: $\mathbf{x}\leftarrow T(\mathbf{x})$
9:        Increment $i$
[…]
11:     Compute $\mathbf{H}=\nabla^{2}p(\mathbf{x})$
12:     if $\max(\text{eigenvalues}(\mathbf{H}))<0$ then
13:        Merge nearby modes within a distance threshold, keeping the highest-probability mode
14:        Add mode to $\mathcal{M}$
15:     end if
16:  end for
17:
18:  Compute Mode Weights: $w_{j}=p(\mathbf{x}_{j})\bigl{|}\!-\!\nabla^{2}\log p(\mathbf{x}_{j})\bigr{|}^{-1/2},\,j=0,\dots,|\mathcal{M}|$
19:  Sample mode: Sample $j\sim\text{Categorical}\left(\{w_{j}\}_{j=0}^{|\mathcal{M}|}\right)$
20:  Return: $\mathbf{x}_{j}$
Appendix C Environment Details
In our experiments, we evaluated Q-FAT in three distinct environments: Kitchen, PushT, and UR3 BlockPush. In the following, we provide a brief description of each environment and its variants.
  • Franka Kitchen: This environment involves a Franka Panda robotic arm with a nine dimensional action space, designed for multi-task manipulation in a simulated kitchen setting (Gupta et al., 2019). The dataset consists of 566 human-collected demonstrations, where each trajectory completes a subset of four out of seven possible tasks in varying orders.
  • PushT: The PushT environment focuses on pushing a T-shaped block to a target position on a table using two dimensional end-effector velocity control (Chi et al., 2023). The dataset includes 206 human-collected demonstrations.
  • UR3 BlockPush: In this task, a UR3 robotic arm is tasked with moving two blocks to designated goal circles on a table (Florence et al., 2022). The dataset exhibits multimodality, as the blocks can be moved in either order. In the unconditional setting, we evaluate whether both blocks reach their respective goals. In the conditional setting, the model is provided with the target positions of the blocks, and performance is assessed based on the order in which the blocks reach their goals.
  • Multimodal Ant: The Multimodal Ant environment (Lee et al., 2024) is a modification of the Ant environment introduced in Schulman et al. (2015) where the goal of the ant is to visit four distinct locations placed on vertices of a square. The multimodality in the dataset arises from the different order the goals are visited. The dataset contains 600 human-collected trajectories
Appendix D Training Hyperparameters
Hyperparameter PushT Image PushT Kitchen UR3 Multimodal Ant
Layers 6 6 6 12 6
Attention heads 8 8 8 8 8
Embedding dimension 128 128 128 128 128
Dropout probability 0.1 0.1 0.1 0.1 0.1
State history size 5 5 10 10 10
Action horizon 5 5 1 1 1
Training epochs 1000 1000 1200 400 1000
Batch size 256 256 128 2048 128
Number of mixtures $k$ 4 4 4 4 4
Maximum learning rate 1e-3 1e-3 1e-3 1e-4 1e-3
Minimum learning rate - - 1e-6 1e-6 1e-6
Learning Rate Schedule - - Cosine Cosine Cosine
Optimizer Adam Adam Adam Adam Adam
Table 2: Environment dependent hyperparameters.